OpenStudy (anonymous):

A line K passes through the points (4, 2) and (-4, 4). A second line h passes through the points (-8, 1) and (8,-3). Is line k parallel to line h? Why or why not?

OpenStudy (anonymous):

find the gradient of the first line = - 1/4 second line = - 1/4 therefore they are parallel

OpenStudy (anonymous):

thanks